# Persistent Homology for Structural Characterization in Disordered Systems

### v2

@@ Abstract
We propose a unified framework based on persistent homology (PH) to characterize both local and global structures in disordered systems. It can simultaneously generate local and global descriptors using the same algorithm and data structure, and has shown to be highly effective and interpretable in predicting particle rearrangements and classifying global phases. Based on this framework, we define a non-parametric metric, the Separation Index (SI), which not only outperforms traditional bond-orientational order parameters in phase classification tasks but also establishes a connection between particle environments and the global phase structure. Our methods provide an effective framework for understanding and analyzing the properties of disordered materials, with broad potential applications in materials science and even wider studies of complex systems.
@@ I Introduction
Topological Data Analysis (TDA) provides a robust and effective tool for representing and analyzing point cloud data, and through persistent homology (PH), a method from algebraic topology [50], we can encode the topological information of point clouds into a vectorizable representation [51, 52, 53]. Based on this approach, we have developed a framework that enables both local and global characterization of disordered systems. This framework can serve as a supplement to traditional physical methods, providing a unified representation that seamlessly interprets how microscopic structures influence macroscopic properties. Besides, it can also generate the high-performance descriptors, enriching the feature engineering or initial structural mapping in downstream tasks of interpretable machine learning.
@@
The second pathway uses a traditional physics approach without ML. We analyzed the topology of both local particle environments and the global system using our proposed new metric, the Separation Index (SI). The SI surpasses the bond-orientation order parameter in classifying the three phases of multi-particle systems, consistent with the finding that high classification accuracy can be achieved with a single variable. By adjusting the neighborhood radius, SI also functions as a mechanistic model, illustrating how long-range order and global symmetry in crystalline materials emerge from the local environment of hard particles. The calculated results of SI align with those of Global Softness, offering a unified interpretation of fluidity and structural order through both ML and traditional physics perspectives.
@@
    Sec. II introduces methods for performing persistent homology analysis on point cloud data (illustrated in Fig. 1 and Fig. 2) and constructing local and global descriptors.
@@
    Sec. III covers machine learning (ML) tasks, including predicting particle rearrangements and phase classification. It details task design, data generation, sample labeling, dataset construction, ML results and feature importance analysis (FIA).
@@
    Sec. IV introduces the Separation Index (SI), a non-parametric metric, designed for classifying both particle types and global phases, and compares it with Global Softness, a metric derived from ML, for prediction of fluidity trends.
  4. 4.
    Sec. V presents results and analysis by comparing four metrics, exploring the relationship between local and global structures (Fig. 4 and Fig. 5), and specifically demonstrating that our SI outperforms traditional order parameters in phase classification (Fig. 5b)) and capturing phase transitions (Fig. 8).
  5. 5.
    Sec. VI concludes the article and discusses the directions of future research.
@@ II Methods
II.1 Vietoris-Rips (VR) Complex
@@ II.5 Traditional Metrics
II.5.1 Mean Square Displacement (MSD)
@@
II.5.2 bond-orientational Order Parameters
@@
III Machine Learning (ML)
III.1 Task Description
@@
III.2 Data Generation
We utilize molecular dynamics (MD) simulations [69, 70] to generate our experimental data. Each simulation records 1000 steps, where one step equals to $100\times(0.2*(m*\sigma^{2}/\epsilon)^{1/2})$ in Lennard-Jones reduced units [71]. The particle count for each system is consistently $\mathscr{N}=864$.
To investigate the potential structural transitions in a isothermal supercooled liquid at different temperatures, we selected a monotonically increasing sequence of 20 temperatures, i.e., $T_{1}$, $T_{2}$, $\cdots$, $T_{20}$. We generate 20 isothermal trajectories of a Lennard-Jones (LJ) system [72] at these temperatures in NPT ensembles. These trajectories are divided into two groups: Traj($\text{Traj}(T^{(1)}_{i})=\text{Traj}(T_{2i-1})$ (Group 1) and Traj($\text{Traj}(T^{(2)}_{i})=\text{Traj}(T_{2i})$ (Group 2), with $i\in[1,10]\cap\mathbb{N}$.
To study the vitrification process, we performed 1000-step linear quenches on two extra Kob-Andersen binary mixtures (KA System) [73, 74], resulting in two extra trajectories $\text{Traj}_{\text{KA}}^{(1)}$ (showing a complete vitrification) and $\text{Traj}_{\text{KA}}^{(2)}$ (remaining liquid).
III.3 Sample Labelling
Identifying Rearrangements. To characterize the tendency of particle rearrangement, we define the positional change $p_{\text{hop},p}(t)$ of a given particle $p$ within a time window $[t-t_{R/2},t+t_{R/2}]$ [75, 76]. If $p_{\text{hop},p}(t)$ exceeds a threshold $p_{c}$ during this interval, the particle is labeled as “soft”; otherwise, it is labeled as “hard” [77]. Specifically, $p_{\text{hop},p}(t)$ is defined by:
$p_{\text{hop},p}(t)=\sqrt{\langle(\mathbb{r}_{p}-\langle\mathbb{r}_{p}\rangle_{B})^{2}\rangle_{A}\cdot\langle(\mathbb{r}_{p}-\langle\mathbb{r}_{p}\rangle_{A})^{2}\rangle_{B}},$ (4)
where $\mathbb{r}_{p}$ is the position vector of particle $p$, and .$\langle.\rangle_{A}$ and .$\langle.\rangle_{B}$ represent averages over the intervals $A=[t-t_{R/2},t]$ and $B=[t,t+t_{R/2}]$, respectively. We set $t_{R/2}=20$ and $p_{c}=0.1$ based on the assumption that most particles in a liquid state are expected to exhibit “soft” behavior. The labels for Task 1 are then derived from these $p_{\text{hop},p}(t)$ values, with the label for particle $p$ in the system at time step $t$ and temperature trajectory $\text{Traj}(T)$ denoted as $y^{(T)}_{p,t}\in\{\text{soft},\text{hard}\}$.
Identifying Phases. The ten Wolde’s order parameter $Q_{6}$ [32] and mean square displacement (MSD) [78] are jointly employed to distinguish phases in disordered systems. In liquids, the MSD shows a linear increase with a significant slope, indicating high particle mobility, whereas in crystals, the MSD remains steady over time, reflecting particle stability. For amorphous materials, the MSD exhibits more complex, intermediate patterns but generally trends closer to the behavior of crystals. Typically, both liquids and amorphous materials exhibit a $Q_{6}$ value near 0 due to their disordered structures, while crystals display higher and more stable $Q_{6}$ values. In Task 2, we use MSD and $Q_{6}$ for phase labeling, denoting the label for trajectory $\text{Traj}(T)$ at time step $t$ as $y^{(T)}_{\text{entire},t}\in\{\text{liquid},\text{crystal},\text{amorphous}\}$. All data are listed in Table 4 of Appendix E for verification.
III.4 Construction of Datasets
@@
III.5 PCA on Feature Matrix
@@
III.6 ML Results and Analysis
We trained our model using Support Vector Machines (SVM) [80], of which the desicion function is denoted as $\mathbb{y}=f(\mathbb{X})=\mathbb{X}\cdot\mathbb{w}+u$, testing both linear [80] and RBF kernels [81]. The implementation of SVM is detailed in Appendix B. The ML results are presented in Table 1. The linear kernel slightly outperformed the RBF kernel in both tasks, but regardless of which kernel was used, the accuracy remained consistently high.
Task 1 (Local) Task 2 (Global)
Models Metrics Variance Variance
98% 50% 98% 45%
SVM (Linear) Accuracy 86.1% 84.2% 96.1% 94.2%
MCC 0.723 0.686 0.763 0.763
SVM (RBF) Accuracy 86.1% 82.4% 95.5% 91.5%
MCC 0.723 0.649 0.760 0.715
Number of Predictors 188 7 122 1
III.7 Feature Importance Analysis (FIA)
III.7.1 The importance of homology classes
$\phi_{0}$ ($H_{0}$) $\phi_{1}$ ($H_{1}$) $\phi_{2}$ ($H_{2}$)
Task 1 (Local) $28.08\%$ $35.29\%$ $36.63\%$
Task 2 (Global) $30.02\%$ $35.05\%$ $34.93\%$
III.7.2 The key feature in phase classification (Task 2)
IV New Metrics
In Sec. III.7.2, we identified a single variable that enables highly accurate phase classification, showing that this variable alone can act as a key factor in distinguishing phases. This finding challenges the traditional reliance on ten Wolde’s order parameter $Q_{6}$, which, although effective at distinguishing between ordered and disordered phases, struggles to differentiate between liquid phases and amorphous phases due to its insensitivity to subtle changes in SRO.
Building on this discovery, we proposed an explicit mapping called the Separation Index (SI, detailed in Eq. 5). In contrast to $Q_{6}$, SI captures the symmetry and order of the system across multiple scales, enabling effective distinction between liquid phases, crystalline phases, and amorphous phases. Importantly, it is derived from physical intuition rather than relying on ML.
Notably, SI is a purely mathematical and non-ML metric derived from the descriptive statistics of persistent homology, providing a complementary perspective to traditional physical methods. For comparison, we also defined a scalar field, Global Softness (detailed in Eq. 6), derived from the ML results, to quantify the overall trend of fluidity in the system over time.
Note that Global Softness is not our original method but an application of the SVM classification hyperplane. It is obtained by averaging the softness of individual particles and is not derived from the unified framework. In contrast, Separation Index (SI) is fully integrated into the framework, serving as a consistent topological metric for both local and global characterizations.
V Results and Discussion
V.1 Bridging the Particle Environments and Entire Systems
V.2 Phase Transitions
We tracked the evolution in four metrics including normalized MSD, $Q_{6}(t)/\mathscr{N}$, $\text{SI}(t)$ and $\mathcal{S}(t)$ over time on ten trajectories in test group (Group 2), i.e., Traj($\text{Traj}(T^{(2)}_{k}),k\in[1,10]\cap\mathbb{N}$, as well as two trajectories in Group 3, i.e., $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$. The data for all metrics are provided in Fig. 17 of Appendix E for reference.
In addition, $\mathcal{S}(t)$ captures the fluidity trend of the system, offering insights into both its current state and potential future behavior. Note that, $\mathcal{S}(t)$ was trained on the training set (Group 1, LJ system) and applied into the test set (Group 2, LJ system) and Group 3 (KA system). When applied to Group 2, a positive $\mathcal{S}(t)$ indicates a trend towards solidification, while a negative value suggests the system is likely to remain fluid. By correlating with potential energy in the system, $\mathcal{S}(t)$ captures both current structural changes and the potential evolution of fluidity over time. As the system transitions from an amorphous phase (see Fig. 8a)) or a liquid phase (see Fig. 8b)) to a crystalline phase, a sharp drop in $\mathcal{S}(t)$ indicates a substantial reduction in fluidity and the stabilization of the crystalline phase. When applied to Group 3 (KA mixture, see Fig. 8c) and Fig. 8d)), $\mathcal{S}(t)$ required a correction to accurately reflect the positive and negative relationships (see Fig. 16 of Appendix D), as the differences in atomic configurations caused systematic shifts in the absolute scale of fluidity. However, it only serves as a reference, and its shift does not impact the presentation of the underlying principles. Therefore, we ignore exploring the correction mechanism as it is not the key factor. If the purpose is solely to track the fluidity on a single trajectory, this correction to align the zero point is not necessary. On the contrary, as mentioned at the beginning of this paragraph, when applied to Group 2 (LJ system), $\mathcal{S}(t)$ required no such adjustment (see Fig. 8b)), indicating that its positive and negative relationships can be directly used without any corrections.
VI Conclusion and Perspective
We employed persistent homology (PH) to develop a unified mathematical framework for both local and global characterization in disordered systems. This approach produces high-performance descriptors for interpretable machine learning (ML) and offers deep insights into the structure-function relationships in these systems. It reveals crucial links between local particle environments and global structures, predicting particle rearrangement and classifying phases in a universal framework. Our proposed Separation Index (SI) effectively captures the differences in order among liquid, amorphous, and crystalline phases in multi-particle systems, providing an explanation for the mechanism of long-range order formation. Besides, SI, in conjunction with Global Softness, achieves precise detection of phase transitions through non-ML and ML pathways, respectively. Furthermore, complexity science suggests that global behavior in a system arises from nonlinear interactions among local components, rather than from simple linear addition. Our unified framework fundamentally avoids the issue of nonlinear adding, while treating local and global characteristics as homogeneous representations. This offers a novel perspective for research in complexity science.
Acknowledgements.
Appendix A The Protocols of MD Simulations
For LJ systems, the parameters of the Lennard-Jones potentials are: $\epsilon=1$, $\sigma=1$, cutoff radius $r_{c}=3.5\sigma$ and mass $m=1M$. We have adopted a tail correction as the truncation scheme. The computational protocol we have followed begins with a linearly quench of the liquid from $T_{\text{init}}=1.25$ to a given temperature $T_{\text{final}}$ in 20 steps. Then, we perform a $1000$ steps equilibration at temperature $T_{\text{final}}$. These simulations have been conducted within the NPT ensemble with an isotropic pressure $P=5.68(\epsilon/{\sigma^{3}})$, enforced via a chain of five thermostats coupled to a Nosé–Hoover barostat [88] (which accounts for the Martyna-Tobias-Klein correction [89]). The damping parameter of the barostat is $0.5t^{*}$, where $t^{*}=0.002(m*\sigma^{2}/\epsilon)^{1/2}$. For convenience, we denote the trajectory corresponding to its temperature $T$ as $\text{Traj}(T)$. All these trajectories were generated using the same initial positions and velocities for each particle. These settings are the same as those adopted in Ref. [90].
For KA mixtures [73, 74], we use the parameter setting: $\epsilon_{AA}=\sigma_{AA}=1$, $\epsilon_{AB}=1.5$, $\sigma_{AB}=0.8$, $\epsilon_{BB}=0.5$, $\sigma_{BB}=0.88$. We started from $T_{\text{init}}=1.25$ and quenched the system to the final temperature $T_{\text{final}}$, which is low enough to guarantee the solidification of the system, either in the crystalline or amorphous phase. The details of these MD simulations are identical to those utilised for the homogeneous LJ system including the system size - 864 particles in total, with the proportion of B particles as $\chi_{B}=0.2$. We generated two trajectories within the time step of $1000$, where one time step equals to $100\times(0.2*(m*\sigma^{2}/\epsilon)^{1/2})$ in LJ reduced units. We denote these two trajectories as $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$ (Group 3). The $T_{\text{final}}$ settings are: (1) for $\text{Traj}_{\text{KA}}^{(1)}$, $T_{\text{final}}=0.5$.; (2) for $\text{Traj}_{\text{KA}}^{(2)}$, $T_{\text{final}}=1.0$.
Appendix B Implementation of SVM
This article uses two kinds of SVMs as $f$:

### v3

@@ Abstract
We propose a unified framework based on persistent homology (PH) to characterize both local and global structures in disordered systems. It can simultaneously generate local and global descriptors using the same algorithm and data structure, and has shown to be highly effective and interpretable in predicting particle rearrangements and classifying global phases. We also demonstrated that using a single variable enables a linear SVM to achieve nearly perfect three-phase classification. Inspired by this discovery, we define a non-parametric metric, the Separation Index (SI), which not only achieves this classification without sacrificing significant performance but also establishes a connection between particle environments and the global phase structure. Our methods provide an effective framework for understanding and analyzing the properties of disordered materials, with broad potential applications in materials science and even wider studies of complex systems.
@@ I Introduction
Topological Data Analysis (TDA) provides a robust and effective tool for representing and analyzing point cloud data, and through persistent homology (PH), a method from algebraic topology [50], we can encode the topological information of point clouds into a vectorizable representation [51, 52, 53]. It is notable that PH has already seen successful applications, such as in studying the global topological characteristics of osmolyte molecular aggregation and hydrogen-bonding networks [54], integrating machine learning to predict the structure-energy relationships of molecular clusters [55], and employing weighted approaches to uncover local topological features and propose new quantitative tools [56].
Based on this approach, we have developed a framework that enables both local and global characterization of disordered systems. This framework can serve as a supplement to traditional physical methods, providing a unified representation that seamlessly interprets how microscopic structures influence macroscopic properties. Besides, it can also generate the high-performance descriptors, enriching the feature engineering or initial structural mapping in downstream tasks of interpretable machine learning.
@@
The second pathway uses a traditional physics approach without ML. We analyzed the topology of both local particle environments and the global system using our proposed new metric, the Separation Index (SI). The SI maintains a comparable level of performance to that achieved by the machine learning model, aligning with the finding that high classification accuracy can be achieved with a single variable. By adjusting the neighborhood radius, SI also functions as a mechanistic model, illustrating how long-range order and global symmetry in crystalline materials emerge from the local environment of hard particles. The calculated results of SI align with those of Global Softness, offering a unified interpretation of fluidity and structural order through both ML and traditional physics perspectives.
@@
    Sec. II introduces the theories and methods involved in this article, including the persistent homology (PH), unified structural characterization, and both ML-based and non-ML approaches.
@@
    Sec. III presents the computational results and provides analysis, including the outcomes and interpretations of both ML-based and non-ML approaches. It also explores phase transitions and their onset by tracking changes in various metrics along the simulated trajectories.
@@
    Sec. IV concludes the article and discusses the directions of future research.
@@ II Methods
II.1 Topological Methods
II.1.1 Vietoris-Rips (VR) Complex
@@ II.2 Traditional Metrics
II.2.1 Mean Square Displacement (MSD)
@@
II.2.2 Bond-orientational Order Parameters
@@
II.2.3 Identifying Rearrangements
To characterize the tendency of particle rearrangement, we define the positional change $p_{\text{hop},p}(t)$ of a given particle $p$ within a time window $[t-t_{R/2},t+t_{R/2}]$ [72, 73]. If $p_{\text{hop},p}(t)$ exceeds a threshold $p_{c}$ during this interval, the particle is labeled as “soft”; otherwise, it is labeled as “hard” [74]. Specifically, $p_{\text{hop},p}(t)$ is defined by:
$p_{\text{hop},p}(t)=\sqrt{\langle(\mathbb{r}_{p}-\langle\mathbb{r}_{p}\rangle_{B})^{2}\rangle_{A}\cdot\langle(\mathbb{r}_{p}-\langle\mathbb{r}_{p}\rangle_{A})^{2}\rangle_{B}}$ (4)
where $\mathbb{r}_{p}$ is the position vector of particle $p$, and .$\langle.\rangle_{A}$ and .$\langle.\rangle_{B}$ represent averages over the intervals $A=[t-t_{R/2},t]$ and $B=[t,t+t_{R/2}]$, respectively.
We set $t_{R/2}=20$ and $p_{c}=0.1$ based on the assumption that most particles in a liquid state are expected to exhibit “soft” behavior. The detailed explanation of threshold selection is provided in the Appendix A.
II.3 Sample Preparation
In sample preparation, molecular dynamics (MD) simulations [75, 76] serve as the core technique. To begin with, we utilized the MD simulations to generate multiple trajectories. Then, we sampled these trajectories to obtain system-level samples (a single frame from a trajectory) and particle-level samples (the local environment of an individual particle within a single frame of the trajectory). Finally, using these samples, we constructed different datasets for both ML and non-ML tasks.
II.3.1 Protocols of MD Simulation
This study involves two MD simulation protocols: Protocol 1 focuses on the isothermal evolution of the LJ system, while Protocol 2 addresses the KA mixture during a linear quenching process. Both protocols utilized a particle count of $\mathscr{N}=864$ to balance the demands and computational efficiency. In the simulation, we use Lennard-Jones reduced units [77], and denote each time step as $100\times(0.2*(m*\sigma^{2}/\epsilon)^{1/2})$.
Protocol 1. For LJ systems, the parameters of the Lennard-Jones potentials are: $\epsilon=1$, $\sigma=1$, cutoff radius $r_{c}=3.5\sigma$ and mass $m=1M$. We have adopted a tail correction as the truncation scheme. The computational protocol we have followed begins with a linearly quench of the liquid from $T_{\text{init}}=1.25$ to a given temperature $T_{\text{final}}$ in 20 steps. Then, we perform a $1000$ steps equilibration at temperature $T_{\text{final}}$. These simulations have been conducted within the NPT ensemble with an isotropic pressure $P=5.68(\epsilon/{\sigma^{3}})$, enforced via a chain of five thermostats coupled to a Nosé–Hoover barostat [78] (which accounts for the Martyna-Tobias-Klein correction [79]). The damping parameter of the barostat is $0.5t^{*}$, where $t^{*}=0.002(m*\sigma^{2}/\epsilon)^{1/2}$. For convenience, we denote the trajectory corresponding to its temperature $T$ as $\text{Traj}(T)$. All these trajectories were generated using the same initial positions and velocities for each particle. These settings are the same as those adopted in Ref. [80], and previously work shown [81] these settings allow us to observe crystal nucleation within a timescale accessible via unbiased MD simulations.
Protocol 2. For KA mixtures [82, 83], we use the parameter setting: $\epsilon_{AA}=\sigma_{AA}=1$, $\epsilon_{AB}=1.5$, $\sigma_{AB}=0.8$, $\epsilon_{BB}=0.5$, $\sigma_{BB}=0.88$. We started from $T_{\text{init}}=1.25$ and quenched the system to the final temperature $T_{\text{final}}$, which is low enough to guarantee the solidification of the system, either in the crystalline or amorphous phase. The details of these MD simulations are identical to those utilised for the homogeneous LJ system including the system size - 864 particles in total, with the proportion of B particles as $\chi_{B}=0.2$. We generated two trajectories within the time step of $1000$, where one time step equals to $100\times(0.2*(m*\sigma^{2}/\epsilon)^{1/2})$ in LJ reduced units.
II.3.2 Group Division
We generated trajectories for three groups according to the protocols provided in Sec. II.3.1. Specifically, Group 1 and Group 2 used Protocol 1, while two trajectories of Group 3 followed Protocol 2.
Group 1 and 2 considered 20 corresponding $T_{\text{final}}$ values which are as follows: $T_{1}=0.5$, $T_{2}=0.53$, $T_{3}=0.55$, $T_{4}=0.57$, $T_{5}=0.6$, $T_{6}=0.63$, $T_{7}=0.65$, $T_{8}=0.67$, $T_{9}=0.7$, $T_{10}=0.73$, $T_{11}=0.75$, $T_{12}=0.77$, $T_{13}=0.8$, $T_{14}=0.83$, $T_{15}=0.85$, $T_{16}=0.88$, $T_{17}=0.9$, $T_{18}=0.93$, $T_{19}=0.95$, $T_{20}=0.97$, i.e, $\text{Traj}(T_{i}),i\in[1,20]\cap\mathbb{N}$. Group $1$ includes temperatures with odd indices, i.e, the trajectories corresponding to $T^{(1)}_{i}=T_{2i-1}$; conversely, Group $2$ includes temperatures with even indices, i.e, the trajectories corresponding to $T^{(2)}_{i}=T_{2i}$, where $i\in[1,10]\cap\mathbb{N}$. Group 3 considered two trajectories, with their corresponding final $T_{\text{final}}$ values being: $T_{\text{final}}=0.5$ and $T_{\text{final}}=1.0$. We denote these two trajectories as $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$, respectively.
In ML tasks, Group 1 was used for training, and Group 2 was used for testing. In non-ML tasks, Groups 1 and 2 were jointly used for descriptive statistics of barcodes. Specifically, the data presented in Table 1, Table 2, Fig. 4, and Fig. 5 were derived from Groups 1 and 2, while Fig. 8a), Fig. 8b), and Fig. 9 were generated by testing the model trained on data from Group 1 with Group 2. The inclusion of Group 3 was solely due to the need for simulating the vitrification of supercooled liquids, which could not be captured from the isothermal evolution trajectories of Groups 1 and 2. Consequently, Fig. 8c) and Fig. 8d) were produced by testing the model trained on data from Group 1 with data from Group 3.
II.3.3 Sample Labelling
Each trajectory is essentially a time series composed of frame-by-frame data, describing the changes in the coordinates of each particle in the system over time. Our study involves two types of samples: particle samples and system samples. A system sample refers to a specific frame in a trajectory, while a particle sample refers to a specific particle within a frame of a trajectory.
The particle samples are labelled according to Eq. 4, which is used to characterize whether the particles will undergo rearrangement within a specific time interval. The labels for particle samples are then derived from these $p_{\text{hop},p}(t)$ values, with the label for particle $p$ in the system at time step $t$ and temperature trajectory $\text{Traj}(T)$ denoted as $y^{(T)}_{p,t}\in\{\text{soft},\text{hard}\}$.
The system samples are jointly labelled by MSD and $Q_{6}$. Firstly, MSD is used to distinguish between liquids and low-mobility solids. In liquids, the MSD shows a linear increase with a significant slope, indicating high particle mobility, whereas in crystals and amorphous solids, the MSD remains steady over time, reflecting particle stability. Note that the MSD of an amorphous material undergoing crystallization will be greater than 0. The difference between two types of solids, amorphous and crystalline, lies in whether long-range order (LRO) is established, and $Q_{6}$ is used as its estimator. Note that both rapid linear quenching of a liquid and incomplete crystallization caused by frustration can lead to the formation of an amorphous state. The former typically suppresses crystallization, resulting in $Q_{6}$ values close to 0, while the latter produces nonzero $Q_{6}$ values that are constant but still significantly smaller than those of crystals. Therefore, we set a threshold of $0.5$ to determine whether the system has crystallized. When $Q_{6}$ exceeds this value, the system is considered to be crystallized.
For system samples, we use MSD and $Q_{6}$ for phase labeling, denoting the label for trajectory $\text{Traj}(T)$ at time step $t$ as $y^{(T)}_{\text{entire},t}\in\{\text{liquid},\text{crystal},\text{amorphous}\}$. All data are listed in Table 4 of Appendix C for verification.
II.3.4 Dataset Construction
II.4 ML-based Approach
II.4.1 Task Description
@@
II.4.2 PCA on Feature Matrix
The feature matrix $\mathbb{X}$ is the result of feature engineering for samples. Specifically, it is constructed by concatenating the vectors $\mathbb{I}^{(P)}_{p}$ and $\mathbb{I}^{(P)}_{\text{entire}}$ corresponding to each sample into a single matrix.
@@ II.4.2 PCA on Feature Matrix
For all ML tasks discussed in this article, the feature matrix undergoes PCA by default, which will not be elaborated further. For simplicity, we will use $\mathbb{X}$ to represent $\mathbb{X}_{\text{PCA}}$ in the following sections.
II.4.3 ML Models
The task of ML is to train a mapping $\mathbb{y}=f(\mathbb{X})$, where $\mathbb{X}$ is the feature matrix mentioned in Sec. II.4.2, and $\mathbb{y}$ is the vector formed by concatenating the labels $y^{(T)}_{p,t}$ or $y^{(T)}_{\text{entire},t}$ for each sample, as described in Sec. II.3.3. In other words, the goal is to train a mapping from the feature matrix to the target properties.
This article uses two kinds of Support Vector Machines (SVM) [85] as $f$: